OpenStudy (anonymous):

A(n) 1090 N man stands in the middle of a frozen pond of radius 6.09 m. He is un- able to get to the other side because of a lack of friction between his shoes and the ice. To overcome this difficulty, he throws his 1.05 kg physics textbook horizontally toward the north shore, at a speed of 4.53 m/s. How long does it take him to reach the south shore?

OpenStudy (anonymous):

This is a momentum problem. Once the man throws the book, the speed he gets multiplied by his mass will be equal to the mass of the boss multiplied by 4.53. So if we set up the momentum equation: (1090/9.81)(his velocity) = (1.05)(4.53) The reason I divided 1090 by 9.81 is because 1090 is the force he is exerting on the ground (his weight). It is his mass multiplied by gravity, so I divided by the acceleration due to gravity. Doing the math, we get roughly 0.043 m/s. From here, it's just basic physics. The man is standing still, so he has no initial acceleration. So we have: distance = velocity * time time = distance/velocity distance = 6.09 m velocity = 0.043 m/s So: time = ~142 s.